OpenStudy (lxelle):

e^-1/2 x ((1+2x)^-1/2 - 1/2(1+2x)^1/2) = 0 find x

OpenStudy (anonymous):

\[\large e^{-\frac{1}{2}x} (1+2x)^{-\frac{1}{2}} - \frac{1}{2}(1+2x)^{\frac{1}{2}} = 0\]

OpenStudy (anonymous):

Is this right?

OpenStudy (lxelle):

i ve edited the ques. sorry my mistake.

OpenStudy (anonymous):

\[\large e^{-\frac{1}{2}x} ( (1+2x)^{-\frac{1}{2}} - \frac{1}{2}(1+2x)^{\frac{1}{2}}) = 0\] Is this good now?

OpenStudy (lxelle):

yeag

OpenStudy (anonymous):

See: \[\large a^{-x} = \frac{1}{(a)^x}\]

OpenStudy (lxelle):

i know how to solve but i couldnt get the answer. could you pls write out the steps for me to get x. thanks.

OpenStudy (anonymous):

So: \[(1 + 2x)^{-\frac{1}{2}} = \frac{1}{(1+2x)^{\frac{1}{2}}}\]

OpenStudy (anonymous):

What answer you got? Tell me that first and allow me to find where you have done mistake(s)..

OpenStudy (lxelle):

x=1/2?

OpenStudy (anonymous):

\[\frac{e^{-x/2}}{(1+2x)^{1/2}}[1 - (1+2x)] = 0 \\ 1 - 1 - 2x = 0 \implies x = 0\]

OpenStudy (lxelle):

what? x=0?

OpenStudy (anonymous):

Nope..

OpenStudy (anonymous):

I missed 1/2 in front..

OpenStudy (lxelle):

huh?

OpenStudy (anonymous):

\[\frac{e^{-x/2}}{(1+2x)^{1/2}}[1 - \frac{1}{2}(1+2x)] = 0 \\ 1 - \frac{1}{2} - x = 0 \implies x = \frac{1}{2}\]

OpenStudy (anonymous):

You are right, why you think you are wrong?

OpenStudy (lxelle):

wait. cant we just solve from ((1+2x)^−1/2−1/2(1+2x)^1/2)=0

OpenStudy (anonymous):

Boxes everywhere..

OpenStudy (lxelle):

huh?

OpenStudy (lxelle):

refresh it.

OpenStudy (anonymous):

I am not able to see what you wrote there, I am seeing boxes.

OpenStudy (anonymous):

Yes you can.. You will get the same..

OpenStudy (lxelle):

how?

OpenStudy (lxelle):

show me the working maybe?

OpenStudy (anonymous):

Just do what I did..

OpenStudy (lxelle):

no. what if i wanna do it that way, the one i ask

OpenStudy (anonymous):

\[\frac{1}{(1+2x)^{1/2}} - \frac{1}{2}(1+2x)^{1/2} = 0 \\ \frac{1}{(1+2x)^{1/2}}[1 - \frac{1}{2}(1+2x)] = 0\]

OpenStudy (anonymous):

Now, put the other term to \(0\), and find \(x\) from there..

OpenStudy (lxelle):

nooo. i mean from (1+2x)^-1/2 = 1/2 (1+2x)^1/2

OpenStudy (lxelle):

how do you solve this.

OpenStudy (anonymous):

Okay..

OpenStudy (anonymous):

\[\frac{1}{(1+2x)^{1/2}} = \frac{1}{2}(1+2x)^{1/2} \]

OpenStudy (lxelle):

uh huh

OpenStudy (anonymous):

Square both the sides, instead of doing "huh", just try to understand what I did..

OpenStudy (lxelle):

i was acknowledging to what you did there.

OpenStudy (anonymous):

Don't square both sides..

OpenStudy (lxelle):

i'm not getting what you trying to say. Youre confusing me

OpenStudy (anonymous):

Multiply by (1+2x)^(1/2) both sides: \[1 = \frac{1}{2}(1+2x) \\ 1 + 2x = 2 \\ 2x = 1 \implies x= \frac{1}{2}\]

OpenStudy (lxelle):

how about 1/2 = (1+2x)^-1/2 / (1+2x) ^1/2

OpenStudy (anonymous):

You can do it in thousand number of ways, I have provided you with enough steps that you can utilize them to understand it..

OpenStudy (lxelle):

no im trying to try out a few steps. thats the only step i got in my mind.

OpenStudy (lxelle):

if you're reluctant to enlighten me please do say so.

OpenStudy (anonymous):

Use the very first formula I gave..

OpenStudy (anonymous):

\[(1+2x)^{-1/2} = ??\]

OpenStudy (lxelle):

isn't in 1 instead of x on the rhs?

OpenStudy (anonymous):

What? Come again..

OpenStudy (lxelle):

1/2 = (1+2x)^-1/2 / (1+2x) ^1/2 this. after solving isn't it 1 instead of x

OpenStudy (anonymous):

Where? Solve it and show..

OpenStudy (lxelle):

(1+2x)^1/2 / (1+2x)^1/2 = 1 no?

OpenStudy (anonymous):

No..

OpenStudy (lxelle):

x?

OpenStudy (anonymous):

\[\frac{1}{2} = \frac{(1+2x)^{-1/2}}{(1+2x)^{1/2}}\]

OpenStudy (anonymous):

Don't do anything else now.. Use that formula which I gave..

OpenStudy (anonymous):

\[(1+2x)^{-1/2} = ??\]

OpenStudy (lxelle):

nevermind lol thanks. youre still not answering my ques. and you insist of using your way.

OpenStudy (anonymous):

nevermind, your question is still not clear to me. Ask it in a good way otherwise show it if I am not getting it..

OpenStudy (anonymous):

And I am stressing my way as I am not getting your way.. :P